OpenStudy (agentnao):

A chemical storeroom has an 80% alcohol solution and a 30% alcohol solution. How many milliliters of each should be used to obtain 50 milliliters of a 60% solution?

OpenStudy (anonymous):

Let x = amount of 80% solution 50-x= amount of 30% solution 50 = number of milliliters of 60% .80x+.30(50-x)=.60*50 .80x+15-.30x=30 .50x+15=30 .50x=15 x=15/.50=30 milliliters of 80% 50-x=20 milliliters of 30%

OpenStudy (agentnao):

I knew that's the formula I needed, but I didn't know what to do for the amounts of the 80% and 30%! Thank you so much.

OpenStudy (anonymous):

no problem